Find f+g, f-g, f/g, f*g, f(g)
f(x)=x squared + 4x-12 g(x)=x-2

Answer by oscargut(891)

Find f+g, f-g, f/g, f*g, f(g)
f(x)=x^2+ 4x-12 g(x)=x-2
f+g = x^2+5x-14
f-g = x^2+3x-10
f*g = (x^2+4x-12)(x-2)=x^3+2x^2-20x+24
f/g = (x^2+4x-12)/(x-2)=(x-2)(x+6)/(x-2)=x+6
f(g)=f(x-2)=(x-2)^2+4(x-2)-12=x^2-4x+4+4x-8-12= x^2-16